OpenStudy (anonymous):

Newton’s law of cooling states that the rate of cooling of an object is proportional to the temperature difference between the object and its surroundings. So the rate of cooling for a bottle of lemonade at a room temperature of 75°F which is placed into a refrigerator with temperature of 38°F can be modeled by dT/dt=k(T-38) where T(t) is the temperature of the lemonade after t minutes and T(0) = 75. After 30 minutes the lemonade has cooled to 60°F, so T(30) = 60. To the nearest degree, what is the temperature of the lemonade after an additional 30 minutes?

OpenStudy (anonymous):

dT/dt=k(T-38) This is a separable differential equation. \[\frac{dT}{dt}=k(T-38) ----->\frac{1}{(T-38)}dT=k*dt\] Integrate both sides: \[\int\limits_{}^{} \frac{1}{(T-38)}dT=\int\limits_{}^{} k*dt\] \[\ln(T-38)=k*t +constant\]

OpenStudy (anonymous):

undo the ln with e

OpenStudy (anonymous):

\[T-38=e^{k*t +constant}\]

OpenStudy (anonymous):

\[T-38=Ae^{k*t}\] A is just the constant pulled out of the exponent

OpenStudy (anonymous):

\[T=38+Ae^{k*t}\]

OpenStudy (anonymous):

T(0) = 75 T(30) = 60

OpenStudy (anonymous):

I don't feel like calculating what A and k are, but it should be doable.

OpenStudy (anonymous):

@Michele_Laino can u help me from here

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we have to solve the differential equation, first. In order to that, I will apply the formula, which I wrote in my tutorial

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

here is my tutorial, for more explanation: http://openstudy.com/study#/updates/552b4f50e4b04e5707c1d380

OpenStudy (anonymous):

@Michele_Laino okay. i will be back in about 35 min. i have to go to lunch rn. but i promise ill be back.

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

@Michele_Laino im back

OpenStudy (michele_laino):

I can rewrite your equation as below: \[\Large \frac{{dT}}{{dt}} - KT = - 38K\] so, comparing that formula, with the general formula of my tutorial, I get: \[\Large \begin{gathered} P\left( t \right) = - K,\quad Q\left( t \right) = - 38K \hfill \\ \hfill \\ {e^{\int P }} = {e^{ - Kt}},\quad {e^{ - \int P }} = {e^{Kt}} \hfill \\ \end{gathered} \] then the solution, is: \[\Large T\left( t \right) = {e^{Kt}}\left( {C + \int { - 38K\;{e^{ - Kt}}dt\;} } \right)\] and finally: \[\Large T\left( t \right) = C{e^{Kt}} + 38\]

OpenStudy (anonymous):

okay so how do I determine the temp after 30 min? @Michele_Laino

OpenStudy (michele_laino):

we have to set T(0)=75, so we get: \[75 = T\left( 0 \right) = C{e^{K \times 0}} + 38 = C + 38\]

OpenStudy (anonymous):

so c=37?

OpenStudy (michele_laino):

yes! Nevertheless, our function T(t) is an increasing function

OpenStudy (michele_laino):

I think there is a typo error in the text of the exercise

OpenStudy (anonymous):

heres the screenshot

OpenStudy (michele_laino):

I think that the rate of cooling, has to be this: \[\frac{{dT}}{{dt}} = - K\left( {T - 38} \right)\] where K>0

OpenStudy (michele_laino):

otherwise, the formula of T above is correct, if K<0

OpenStudy (michele_laino):

Let's see if that conjecture is true

OpenStudy (michele_laino):

we have: \[T\left( t \right) = 37{e^{Kt}} + 38\] now I substitute the condition: T(30)=60, and I can write: \[60 = T\left( {30} \right) = 37{e^{30K}} + 38\] and: \[{e^{30K}} = \frac{{60 - 38}}{{37}} = ...?\]

OpenStudy (michele_laino):

@familyguymath

OpenStudy (anonymous):

0.59

OpenStudy (michele_laino):

ok! better is 22/37

OpenStudy (michele_laino):

so we have: \[\Large {e^{30K}} = \frac{{22}}{{37}}\]

OpenStudy (anonymous):

right

OpenStudy (michele_laino):

now, we have to compute this: T(30+30)=...?

OpenStudy (michele_laino):

namely: T(60)=...?

OpenStudy (michele_laino):

we can write: \[\Large \begin{gathered} T\left( {60} \right) = 37{e^{60K}} + 38 = 37{\left( {{e^{30K}}} \right)^2} + 38 = \hfill \\ \hfill \\ = 37 \times {\left( {\frac{{22}}{{37}}} \right)^2} + 38 = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

@familyguymath

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

51.08

OpenStudy (anonymous):

or 1890/37

OpenStudy (michele_laino):

ok! That's right!

OpenStudy (michele_laino):

T(60)=51.08 °F

OpenStudy (anonymous):

alright so what about the second half?

OpenStudy (michele_laino):

we have to solve this equation: \[\Large 55 = 37{e^{K{t_0}}} + 38\] and we have to compute t_0

OpenStudy (michele_laino):

so, we can write: \[\Large {e^{K{t_0}}} = \frac{{55 - 38}}{{37}} = ...?\]

OpenStudy (anonymous):

17/37

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, I can write this: \[\Large {e^{\left( {30K} \right) \times \left( {\frac{{{t_0}}}{{30}}} \right)}} = {\left( {{e^{30K}}} \right)^{{t_0}/30}} = {\left( {\frac{{22}}{{37}}} \right)^{{t_0}/30}} = \frac{{17}}{{37}}\]

OpenStudy (michele_laino):

so: \[\Large {t_0} = \frac{{\log \left( {17/37} \right)}}{{\log \left( {22/37} \right)}} = \frac{{\log 17 - \log 37}}{{\log 22 - \log 37}} = ...?\]

OpenStudy (michele_laino):

@familyguymath

OpenStudy (anonymous):

1.49594

OpenStudy (michele_laino):

yes! Nevertheless I compute another time t_0

OpenStudy (michele_laino):

I try to check my computation

OpenStudy (anonymous):

okay and?

OpenStudy (michele_laino):

sorry, the right formula, is: \[\Large {t_0} = 30 \times \frac{{\log \left( {17/37} \right)}}{{\log \left( {22/37} \right)}} = 30 \times \frac{{\log 17 - \log 37}}{{\log 22 - \log 37}} = ...?\]

OpenStudy (michele_laino):

@familyguymath

OpenStudy (anonymous):

45 @Michele_Laino

OpenStudy (michele_laino):

that's right!